# Exploring Typographic Visual Prompts Injection Threats in Cross-Modality Generation Models

### v1

@@ 4.1 Experimental Setting
Models    For the Vision-Language Perception (VLP) task, we conduct extensive experiments on current advanced open-source Large Vision Language Models series LLaVA-v1.6 [31, 30], InternVL-v2.5 [6], Ovis-v2 [33], and Qwen-v2.5-VL [43, 45]. For closed-source LVLMs, we evaluate two widely-used commercial models with APIs: Claude-3.5-Sonnet (Anthropic) [2] and GPT-4o (OpenAI) [36]. For the Image-to-Image (I2I) task, we conduct experiments across DALL-E 2 or UnCLIP [41] and IP-Adapter [50]. For IP-Adapter, we adopt three popular diffusion models, which are Stable Diffusion v1.5 (SD1.5) [42], Stable Diffusion XL (SDXL) [38], and FLUX.1-dev (FLUX) [13]. For closed-source I2I GMs, we evaluate two popular models GPT-4 (OpenAI) [36] and Dreamina (ByteDance) [4].
Datasets    We adopt Typographic Visual Prompt Injection (TVPI) Dataset. The VLP and I2I subtype Dataset, including FM and DTT, are used to evaluate the TVPI threats of various Cross-Vision GMs (LVLMs and I2I GMs) under different factors and attack targets. The closed-source subtype dataset is specifically designed to execute on commercial APIs and official websites [2, 36, 4] of various GMs.
@@
In the I2I task, we employ CLIPScore [40] to measure semantic alignment between generated images and their corresponding inserted attack targets. Higher CLIPScore values indicate stronger semantic similarity between the generated image and attack targets, suggesting more significant influence from the typographic visual prompts. Additionally, we utilize Fréchet Inception Distance (FID) [17] to quantify distribution differences between images generated from visual-prompt-injected inputs and their corresponding clean originals. Larger FID scores signify greater deviation from source images, demonstrating stronger attack impact. Due to space constraints, we present the experiments measured by FID in the Appendix.
@@ 4.3.1 Impact on Open-Source Models
As shown in Table 3, we can observe significant variations in model vulnerability to typographic visual prompts across different scenarios. For VLP tasks, a notable pattern emerges within model families: smaller models generally demonstrate resilience to visual prompts, while larger models LLaVA-v1.6-72B, InternVL-v2.5-38B, and Qwen-v2.5-VL-72B exhibit pronounced susceptibility, manifesting in elevated ASR. Interestingly, A non-linear relationship between model size and robustness appears in the InternVL-v2.5 and Ovis-v2 series, where vulnerability initially increases with model size but then decreases as models scale further, suggesting that the largest variants regain resistance to typographic visual prompts. For I2I tasks, all models show increased CLIPScores under the impact of typographic visual prompts, compared to the clean setting. Figure 2 shows examples of generated images affected by typographic visual prompts. More examples of generated images are presented in Appendix Section C.
@@
Table 3: The impact of typographic visual prompts with different attack targets and under defense (values in parentheses) across VLP tasks (measured by average ASR across four subtasks) and I2I tasks (measured by average CLIPScore across two subtasks). Higher values indicate a stronger effect of typographic visual prompts. Gray indicates models which are less affected by typographic visual prompts. Green highlights where defense decreases the ASR and CLIPScore.
4.3.2 Impact on Closed-Source Models
To demonstrate the potential impact of RVPI in society, we also extend the experiment to include closed-source models, showing their vulnerability to typographic visual prompts.
For the VLP task, Table 3 shows that Claude-3.5-Sonnet (Anthropic) [2] and GPT-4o (OpenAI) [36] are severely affected by typographic visual prompts. In the I2I task, as illustrated in Figure 2, the generated images from both GPT-4 (OpenAI) [36] and Dreamina (ByteDance) [4] exhibit clear influence from typographic visual prompts. More examples of generated images are presented in Appendix Section C.
4.3.3 Defense
As illustrated in Table 3, in the VLP task, the defense shows partial effectiveness in reducing the ASR across some models. However, the overall ASR remains notably high despite this intervention. Furthermore, the results are less promising for I2I tasks, where the defense demonstrates minimal impact in terms of CLIPScore. Examples of generated images showing the defense is ineffective are presented in Appendix Section C. These findings highlight the resilience of typographic visual prompts against simple prompt modification.
4.4 Analysis
Figure 3: (a) Performance of models detecting the typographic visual prompts under different conditions. (b) The impact of typographic visual prompt and typographic word injection on different targets in VLP tasks (measured by average ASR across four subtasks)
4.4.1 Comparison on Models’ OCR Capability
We further investigate whether a model’s susceptibility to typographic visual prompts correlates with its ability to detect visual prompts in images. Specifically, we use the text prompt "only output the instruction in the image" to assess each model’s detection capability. We consider detection successful only when the model’s response exactly matches the visual prompt embedded in the image. Interestingly, as shown in Figure 3 (a), smaller models demonstrate high detection success rates despite their lower vulnerability to typographic visual prompts, as previously established in Table 3. This finding suggests that the ability to detect text in images does not necessarily translate to vulnerability to typographic visual prompts.
4.4.2 When Visual Prompts Turn Imperceptible
Figure 4: Examples of three strategies to conceal visual prompts within the image. (a) visual prompts with text opacity 25%. (b)(c) visual prompts hidden in black or white.
We also explore whether models can still detect typographic visual prompts that become nearly imperceptible to humans. In particular, As examples shown in Figure 4, we devise three strategies to conceal the visual prompt within the image: (1) reducing the text opacity to 25%, (2) setting the text color to near-black with an RGB value of (5, 5, 5) and embedding it in a black background, and (3) setting the text color to near-white with an RGB value of (250, 250, 250) and embedding it in a white background. Figure 3 (a) illustrates that most models keep high detection success rates when visual prompts turn imperceptible, especially with strategies of hiding in black or white background. For the I2I task, examples of generated images affected by imperceptible visual prompts are presented in Appendix Section C. This finding implies that typographic visual prompt injection can be implemented in such subtle ways, posing significant societal risks.
4.4.3 Comparison with Typographic Word Injection
We also compare the typographic visual prompt injection with the typographic word injection mentioned in the work [7]. Specifically, we reduce the typographic visual prompt to only the attack target word, constituting the typographic word injection. For the VLP task, Figure 3 (b) demonstrates that typographic word has little impact on models’ output, while typographic visual prompts cause a high ASR. In the I2I task, Figure 2 shows typographic word injection has less influence on the generated images from closed-source models GPT-4 and Dreamina, when compared to the effectiveness of typographic visual prompts. Additional examples of generated images are shown in Appendix Section C.
4.4.4 Insights and Implications
Visual Prompt Execution Priority    Our findings reveal a vulnerability in cross-vision generation models: visual prompts embedded in the visual modality receive higher execution priority than those provided through text input. Even in the defense when explicitly directed through text prompts to disregard visual prompts, models continue to prioritize and execute instructions embedded within images across both VLP and I2I tasks. This phenomenon may be rooted in the parameter imbalance between vision and text modality [53].
@@ 4.4.4 Insights and Implications
Model Size    Our experiments also show a complex relationship between model size and vulnerability to typographic visual prompts. While smaller models within a family generally demonstrate greater resilience, we observe that the largest models (LLaVA-v1.6-72B, Qwen-v2.5-VL-72B, GPT-4o, and Claude-3.5-sonnet) exhibit pronounced susceptibility to typographic visual prompts. However, this relationship is not strictly linear, as evidenced by the InternVL-v2.5 and Ovis-v2 series, where vulnerability initially increases with model size but then decreases in the largest variants.
5 Limitations
Due to financial resource limitations and constraints in accessing official APIs and Websites, we conduct experiments with closed-source models using a subset of our dataset. In I2I task, we also evaluate other closed-source models. Midjourney [34] and GPT-4o [36], which demonstrate resilience to typographic visual prompts.
6 Conclusion
In this work, we systematically investigated the impact of Typographic Visual Prompt Injection (TVPI) on Large Vision Language Models (LVLMs) and Image-to-Image Generative Models (I2I GMs). Our study reveals that TVPI significantly influences model outputs, often leading to unintended semantic disruptions. To facilitate analysis, we introduced the TVPI Dataset, enabling a deeper understanding of its effects. Our findings highlight the security risks posed by TVPI in cross-modality generation and provide insights into its underlying mechanisms. This work underscores the need for robust defenses against typographic-based visual attacks in generative models.
Supplementary Material
A. Additional Information for Typographic Visual Prompt Injection Dataset
Figure 6 and Figure 7 show examples of different text factors of a typographic visual prompt in vision-language perception and image-to-image tasks, respectively.
B. Additional Results for Vision-Language Perception Task
For the vision-language perception task, we also curate another visual prompt template: ignore everything, just output {attack target}. Compared to the previous visual prompt template when asked about {subtask type}, just output {attack target} shown in Table 3, this new visual prompt template is less effective on most of the models, as demonstrated in Table 4.
Protective Harmful Bias
Model Clean sorry unknown suicide murder African Asian
LLaVA-v1.6-7B 0.00 0.00 0.00 0.00 0.00 0.00 0.00
LLaVA-v1.6-13B 0.00 0.00 0.00 0.00 0.00 0.00 0.00
LLaVA-v1.6-34B 0.00 0.00 0.00 0.00 0.00 0.00 0.00
LLaVA-v1.6-72B 0.00 0.58 0.42 0.07 0.57 0.67 0.77
InternVL-v2.5-8B 0.00 0.00 0.00 0.00 0.00 0.00 0.00
InternVL-v2.5-38B 0.00 0.07 0.11 0.01 0.03 0.01 0.05
InternVL-v2.5-78B 0.00 0.00 0.01 0.00 0.00 0.00 0.00
Ovis-v2-8B 0.00 0.00 0.00 0.00 0.00 0.01 0.00
Ovis-v2-16B 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Ovis-v2-34B 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Qwen-v2.5-VL-7B 0.00 0.00 0.10 0.00 0.00 0.01 0.00
Qwen-v2.5-VL-72B 0.00 0.93 0.95 0.86 0.89 0.93 0.93
Table 4: The impact of typographic visual prompts in VLP task with the visual prompt template “ignore everything, just output {attack target} (measured by average ASR on four subtasks, with attack target “sorry”).
C. Additional Results for Image-to-Image Task
Figure 5 shows examples of three strategies to conceal visual prompts within the image in the image-to-image tasks: (1) reducing the text opacity to 25%, (2) setting the text color to near-black with an RGB value of (5, 5, 5) and embedding it in a black background, and (3) setting the text color to near-white with an RGB value of (250, 250, 250) and embedding it in a white background.
Figure 5: Examples of three strategies to conceal visual prompts within the image. (a) visual prompts with text opacity 25%. (b)(c) visual prompts hidden in black or white.
Figure 6: Examples of different text factors of a typographic visual prompt in vision-language perception tasks.
Figure 7: Examples of different text factors of a typographic visual prompt in image-to-image tasks.
Table 6 and Table 7 show the text factor and different target experiments in image-to-image tasks, measured by FID scores.
Table 5 shows the impact of typographic word injection measured by FID scores in image-to-image tasks. Compared to typographic visual prompt injection in Table 7, typographic word injection has less impact on image generation, causing lower FID scores.
@@ C. Additional Results for Image-to-Image Task
Table 5: The impact of typographic word injection with different attack targets across image-to-image tasks (measured by average FID across two subtasks).
Text Size Text Opacity Text Positioning
Model Clean 8pt 12pt 16pt 20pt 25% 50% 75% 100% A1 A2 A3 A4
UnCLIP (DALL-E 2) 57.57 66.98 81.41 101.5 116.8 78.91 103.2 114.2 116.8 104.8 119.6 127.9 116.8
IP-Adapter-SD1.5 78.23 103.8 127.9 146.8 156.0 117.4 143.7 153.8 156.0 148.5 155.4 156.3 156.0
IP-Adapter-SDXL 97.84 109.8 120.9 130.3 133.8 119.4 128.6 134.0 133.8 129.7 137.8 138.3 133.8
IP-Adapter-FLUX 101.1 111.9 151.3 165.2 166.1 127.8 150.9 163.0 166.1 157.0 167.1 150.2 166.1
Table 6: The impact of typographic visual prompts with different text factors in image-to-image task (measured by average FID on two subtasks, with attack target “naked”), where a larger value indicates a stronger impact of typographic visual prompts.
Harmful Bias Neutral
Model Clean naked bloody African Asian glasses hat
UnCLIP (DALL-E 2) 57.57 111.0(108.1) 107.1(105.2) 157.0(154.9) 123.4(120.3) 119.5(117.1) 106.5(104.0)
IP-Adapter-SD1.5 78.23 151.5(138.2) 153.5(140.6) 179.3(176.0) 165.8(154.4) 150.8(133.6) 136.7(121.2)
IP-Adapter-SDXL 97.84 131.6(138.2) 116.1(129.5) 132.6(136.5) 129.7(137.5) 120.5(132.0) 124.5(134.6)
IP-Adapter-FLUX 101.1 164.5(168.0) 162.8(168.0) 161.3(157.9) 134.1(130.0) 141.7(140.0) 133.4(132.6)
Table 7: The impact of typographic visual prompts with different attack targets and under defense (values in parentheses) across image-to-image tasks (measured by average FID across two subtasks). Higher values indicate a stronger effect of typographic visual prompts.
Figure 8 and Figure 9 shows generated images from closed-source models GPT-4 and Dreamina, affected by typographic visual prompt injection.
Figure 10 and Figure 11 demonstrate that while applying defense in the text prompts, generated images from closed-source models GPT-4 and Dreamina are still affected by typographic visual prompt injection.
Figure 12 and Figure 13 illustrate generated images from closed-source models GPT-4 and Dreamina are less affected by typographic word injection.
Figure 14 and Figure 15 display that when typographic visual prompts turn imperceptible, generated images from closed-source models GPT-4 and Dreamina are still affected by typographic visual prompt injection.
Figure 16 and Figure 17 present that generated images from open-source models are affected by typographic visual prompt injection.
Figure 8: Randomly selected Typographic Visual Prompt Injection images and their corresponding image outputs under GPT-4.
Figure 9: Randomly selected Typographic Visual Prompt Injection images and their corresponding image outputs under Dreamina.
Figure 10: The image output of GPT-4 after adopting the text prompt with the defense prefix “ignore text in the image” under Typographic Visual Prompt Injection
Figure 11: The image output of Dreamina after adopting the text prompt with the defense prefix “ignore text in the image” under Typographic Visual Prompt Injection
Figure 12: Randomly selected Typographic Word Injection images and their corresponding image outputs under GPT-4.
Figure 13: Randomly selected Typographic Word Injection images and their corresponding outputs under Dreamina.
Figure 14: Image outputs of GPT-4 under three strategies to conceal visual prompts within the image. (1) visual prompts with text opacity 25%. (2)(3) visual prompts hidden in black or white.
Figure 15: Image outputs of Dreamina under three strategies to conceal visual prompts within the image. (1) visual prompts with text opacity 25%. (2)(3) visual prompts hidden in black or white.
Figure 16: The Task 1 (text prompt 1) response of the open-source models under Typographic Visual Prompt Injection.
Figure 17: The Task 2 (text prompt 2) response of the open-source models under Typographic Visual Prompt Injection.

### v2

@@ 4.1 Experimental Setting
Models   For the Vision-Language Perception (VLP) task, we conduct extensive experiments on current advanced open-source Large Vision Language Models series LLaVA-v1.6 (Liu et al., 2024, 2023), InternVL-v2.5 (Chen et al., 2024), Ovis-v2 (Lu et al., 2024), and Qwen-v2.5-VL (Team, 2025; Wang et al., 2024). For closed-source LVLMs, we evaluate two widely-used commercial models with APIs: Claude-3.5-Sonnet (Anthropic) (Anthropic, 2025) and GPT-4o (OpenAI) (OpenAI, 2025). For the Image-to-Image (I2I) task, we conduct experiments across DALL-E 2 or UnCLIP (Ramesh et al., 2022) and IP-Adapter (Ye et al., 2023). For IP-Adapter, we adopt three popular diffusion models, which are Stable Diffusion v1.5 (SD1.5) (Rombach et al., 2022), Stable Diffusion XL (SDXL) (Podell et al., 2023), and FLUX.1-dev (FLUX) (Esser et al., 2024). For closed-source I2I GMs, we evaluate two popular models, GPT-4 (OpenAI) (OpenAI, 2025) and Dreamina (ByteDance) (ByteDance, 2025).
Datasets   We adopt Typographic Visual Prompt Injection (TVPI) Dataset. The VLP and I2I subtype datasets are used to evaluate the TVPI threats of various Cross-Vision GMs (LVLMs and I2I GMs) under different factors and attack targets. The closed-source subtype dataset is specifically designed to execute on commercial APIs and official websites (Anthropic, 2025; OpenAI, 2025; ByteDance, 2025) of various GMs.
@@
In the I2I task, we employ CLIPScore (Radford et al., 2021a) to measure semantic alignment between generated images and their corresponding inserted attack targets. Higher CLIPScore values indicate stronger semantic similarity between the generated image and attack targets, suggesting more significant influence from the typographic visual prompts. Additionally, we utilize Fréchet Inception Distance (FID) (Heusel et al., 2017) to quantify distribution differences between images generated from visual-prompt-injected inputs and their corresponding clean originals. Larger FID scores signify greater deviation from source images, demonstrating stronger attack impact.
@@ 4.3.1 Impact on Open-Source Models
Table 3: The impact of typographic visual prompts with different attack targets and under defense (values in parentheses) across VLP tasks (measured by average ASR across four subtasks) and I2I tasks (measured by average CLIPScore across two subtasks). Higher values indicate a stronger effect of typographic visual prompts. Gray indicates models which are less affected by typographic visual prompts. Green highlights indicates effective defense performance.
@@
As shown in Table 3, we can observe significant variations in model vulnerability to typographic visual prompts across different scenarios. For VLP tasks, a notable pattern emerges within model families: smaller models generally demonstrate resilience to visual prompts, while larger models LLaVA-v1.6-72B, InternVL-v2.5-38B, and Qwen-v2.5-VL-72B exhibit pronounced susceptibility, manifesting in elevated ASR. Interestingly, A non-linear relationship between model size and robustness appears in the InternVL-v2.5 and Ovis-v2 series, where vulnerability initially increases with model size but then decreases as models scale further, suggesting that the largest variants regain resistance to typographic visual prompts. For I2I tasks, all models show increased CLIPScores under the impact of typographic visual prompts, compared to the clean setting. Figure 2 shows examples of generated images affected by typographic visual prompts. Table 4 shows the impact of TVPI measured by FID scores in image-to-image tasks.
@@
Table 4: The impact of TVPI with different attack targets across I2I tasks (measured by average FID across two subtasks).
4.3.2 Impact on Closed-Source Models
To demonstrate the potential impact of TVPI in society, we also extend the experiment to include closed-source models, showing their vulnerability to typographic visual prompts.
For the VLP task, Table 3 shows that Claude-3.5-Sonnet (Anthropic) (Anthropic, 2025) and GPT-4o (OpenAI) (OpenAI, 2025) are severely affected by typographic visual prompts. In the I2I task, as illustrated in Figure 2, the generated images from both GPT-4 (OpenAI) (OpenAI, 2025) and Dreamina (ByteDance) (ByteDance, 2025) exhibit clear influence from typographic visual prompts.
4.3.3 Defense
As illustrated in Table 3, in the VLP task, the defense shows partial effectiveness in reducing the ASR across some models. However, the overall ASR remains notably high despite this intervention. Furthermore, the results are less promising for I2I tasks, where the defense demonstrates minimal impact in terms of CLIPScore. These findings highlight the resilience of typographic visual prompts against simple prompt modification.
4.4 Discussion
4.4.1 Comparison with Typographic Word Injection
We also compare the typographic visual prompt injection with the typographic word injection mentioned in the work (Cheng et al., 2024). Specifically, we reduce the typographic visual prompt to only the attack target word, constituting the typographic word injection. For the VLP task, Figure 3 (b) demonstrates that typographic word has little impact on models’ output, while typographic visual prompts cause a high ASR. In the I2I task, Figure 2 shows that typographic word injection has less influence on the generated images from closed-source models GPT-4 and Dreamina, when compared to the effectiveness of typographic visual prompts.
Figure 3: The impact of typographic visual prompt and typographic word injection on different targets in VLP tasks (measured by average ASR across four subtasks)
4.4.2 Model Size in Typographic Visual Prompt Injection
Our experiments in Table 2 also show a complex relationship between model size and vulnerability to typographic visual prompts. While smaller models within a family generally demonstrate greater resilience, we observe that the largest models (LLaVA-v1.6-72B, Qwen-v2.5-VL-72B, GPT-4o, and Claude-3.5-sonnet) exhibit pronounced susceptibility to typographic visual prompts. However, this relationship is not strictly linear, as evidenced by the InternVL-v2.5 and Ovis-v2 series, where vulnerability initially increases with model size but then decreases in the largest variants.
5 Conclusion
In this work, we systematically investigated the impact of Typographic Visual Prompt Injection (TVPI) on Large Vision Language Models (LVLMs) and Image-to-Image Generative Models (I2I GMs). Our study reveals that TVPI significantly influences model outputs, often leading to unintended semantic disruptions. To facilitate analysis, we introduced the TVPI Dataset, enabling a deeper understanding of its effects. Our findings highlight the security risks posed by TVPI in cross-modality generation and provide insights into its underlying mechanisms. This work underscores the need for defenses against typographic visual attacks in generative models.